# On the spectrum of Sturmian Hamiltonians of bounded type in a small coupling regime

### v1

@@ 1 Introduction
This type of potential gives a self-adjoint operator that is directly associated to a mathematical model of quasicrystals. A survey of results concerning these types of operators, especially Schrödinger operators, can be found in [DEG].
We focus our attention to the case when the frequency has only finitely many different terms that appear in its continued fraction expansion (i.e. of bounded-type) and the coupling is weak (i.e. $\lambda$ is near 00), and investigate the dimension of the spectrum of $H_{\lambda,\alpha,\omega}$ in this context. Since $R_{\alpha}$ is minimal, it is well-known (see Theorem 4.9 in [DF]) that the spectrum of $H_{\lambda,\alpha,\omega}$ is independent of $\omega$, so we will denote it by $\sigma_{\lambda,\alpha}$.
@@
In the case that $\alpha=\frac{\sqrt{5}-1}{2}$, the operator $H_{\lambda,\alpha,\omega}$ is known as the Fibonacci Hamiltonian. In this case, many questions regarding the box-counting and Hausdorff dimensions of the spectrum have been answered (see [D1, D2, DEG, DGY, Su3] for surveys).
@@
For weak coupling constant, many dimension results about the spectrum have been obtained through the analysis of trace map formalism. There is a fundamental link between the spectrum $\sigma_{\lambda,\alpha}$ and the dynamics of the so-called trace maps over a family of cubic surfaces. More precisely, for each $a\in\mathbb{N}$, there is a corresponding trace map $T_{a}$ which leaves invariant a corresponding family of cubic surfaces $\{S_{\lambda}\}_{\lambda>0}$, and the dynamics of the trace maps over $S_{\lambda}$ in the prescribed order that comes exactly from the continued fraction expansion of $\alpha$, gives rich information about the spectrum $\sigma_{\lambda,\alpha}$ (see [Su1, Ra]). For example, the continued fraction expansion of $\alpha=\frac{\sqrt{5}-1}{2}$ contains only 1’s, so here, only the Fibonacci trace map $T_{1}$ plays an important role. Many authors have examined dynamical properties of this trace map (see [BGJ, BR, C, Ca, DG1, DG2, DG3, DG4, DG5, KKT]), where questions concerning hyperbolicity, Markov partitions, entropy, and more have been addressed. It turns out that the non-wandering set of the Fibonacci trace map restricted to any $S_{\lambda}$, $\lambda>0$, is a locally maximal hyperbolic set that is homeomorphic to a Cantor set. This was proved for $\lambda\geq 16$ in [C], for small $\lambda$ in [DG1], and for all $\lambda$ in [Ca]. When $\alpha$ is eventually periodic, one needs to investigate the non-wandering set of a hyperbolic map over $S_{\lambda}$, where this map is precisely a composition of finitely many trace maps, in which case several well-developed techniques from hyperbolic dynamics apply.
In this paper, to work with the weakened assumption that $\alpha$ is of bounded-type, we incorporate non-stationary versions of various dynamical notions such as stable manifolds, Markov rectangles, and a bounded distortion estimate from [BM].
@@ 1.1 Background Material
A main ingredient for achieving such results about the spectrum $\sigma_{\lambda,\alpha}$ is to investigate the dynamics of the so-called trace maps.
1.1.1 Trace Maps
Definition (Trace Maps).
We define the trace maps to be the family of maps given by
$T_{a}:\mathbb{R}^{3}\to\mathbb{R}^{3},\quad\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}\mapsto\begin{pmatrix}x_{1}U_{a}(x_{2})-x_{3}U_{a-1}(x_{2})\\ x_{1}U_{a-1}(x_{2})-x_{3}U_{a-2}(x_{2})\\ x_{2}\end{pmatrix},$
where $U_{a}(t)$ are the Chebyshev polynomials of the second kind. That is,
$U_{-1}(t)=0,\ U_{0}(t)=1,\ U_{a+1}(t)=2tU_{a}(t)-U_{a-1}(t).$
For example, the first two trace maps are given by
$T_{1}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}2x_{1}x_{2}-x_{3}\\ x_{1}\\ x_{2}\end{pmatrix}$
and
$T_{2}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}4x_{1}x_{2}^{2}-2x_{2}x_{3}-x_{1}\\ 2x_{1}x_{2}-x_{3}\\ x_{2}\end{pmatrix}.$
It is a calculation to check that for the maps
$G:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3},\ \text{via}\ G\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}2x_{1}x_{3}-x_{2}\\ x_{1}\\ x_{3}\end{pmatrix}$
and
$H:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3},\ \text{via}\ H\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}x_{1}\\ x_{3}\\ x_{2}\end{pmatrix},$
we have
$T_{1}=G\circ H,\ T_{a}=G^{a-1}\circ T_{1}=G\circ T_{a-1}=G^{a}\circ H.$ (1.2)
@@ 1.1.1 Trace Maps
Each surface $S_{\lambda}$ is invariant under each trace map and from here, we will consider the restriction of each $T_{a}$ to $S_{\lambda}$. Given a sequence $\overline{a}=(a_{n})$ in $\mathbb{N}$, denote
$\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right):=\left\{\boldsymbol{x}\in S_{\lambda}:\left(T_{a_{n}}\circ\cdots T_{a_{1}}(\boldsymbol{x})\right)_{n\in\mathbb{N}}\ \text{is bounded}\right\}.$
@@
which is contained in $S_{\lambda}$, then our goal is to estimate the Hausdorff dimension of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$ as $\lambda\rightarrow 0$. We will estimate, from below, the Hausdorff dimension of a subset of $\Lambda_{\lambda}^{\text{bnd}}$ that consists of all points that remain in a central portion of the surface $S_{\lambda}$, under our prescribed iteration of trace maps. We will now go through some preliminaries and then outline the structure of the paper in Section 1.2.
1.1.2 Contraction Mappings
Proposition 1.
$d\left(f_{n}(x),f_{n}(y)\right)\leq\mu d(x,y),$
for all $x,y\in K$. Then, the following hold:
  • (1)
    There is a $x^{*}\in K$ such that
    for any $x\in K$.
  • (2)
    $d\left(h_{n}(x),h_{n}(y)\right)\leq\mu^{\prime}d(x,y)\ \text{and}\ \sup_{n}d_{C^{0}(K)}(g_{n},h_{n})<\epsilon^{\prime},$
    for some $\mu^{\prime}\in(0,1)$ and for all $x,y\in K$, then
    $d\left(x^{*},y^{*}\right)<\epsilon,$
    where $y^{*}$ is the element such that
    for each $x\in X$.
Proof.
  • (1)
    By compactness of $K$, there is an $M>0$ so that $d(x,y)\leq M$ for all $x,y\in M$. Then, by induction on $n$, we obtain that for any $x\in K$,
@@ 1.1.2 Contraction Mappings
    $d\left(f_{1}\circ\cdots\circ f_{n}(x),f_{1}\circ\cdots\circ f_{m}(x)\right)\leq\mu^{n}d\left(x,f_{n+1}\circ\cdots\circ f_{m}(x)\right)\leq\mu^{n}M,$
    for every $n$ and $m$ such that $n\leq m$, so that $\left(f_{1}\circ\cdots\circ f_{n}(x)\right)$ is a Cauchy sequence. By the completeness of $X$ and compactness of $K$, there must be a $x^{*}\in K$ so that $\lim\limits_{n\rightarrow\infty}f_{1}\circ\cdots\circ f_{n}(x)=x^{*}.$
    Now, let $x^{\prime}\in K$. Then, again by induction on $n$, we have
    $d\left(f_{1}\circ\cdots f_{n}(x),f_{1}\circ\cdots\circ f_{n}\left(x^{\prime}\right)\right)\leq\mu^{n}d\left(x,x^{\prime}\right)\leq\mu^{n}M,$
    for any $n$, and hence $\lim\limits_{n\rightarrow\infty}f_{1}\circ\cdots\circ f_{n}\left(x^{\prime}\right)=x^{*}.$
  • (2)
    for any $x\in K$. We can deduce that
1.1.3 Hyperbolic Dynamics
Suppose that $M$ is a smooth compact Riemannian surface without boundary. For a $C^{r}$, $r\geq 1$, diffeomorphism $f:M\rightarrow M$, an $f$-invariant set $\Lambda\subset M$ is a hyperbolic set for $f$ provided there are values $0<c<1<\mu$ such that for each $x\in\Lambda$, there exist subspaces $H_{x}$ and $V_{x}$, that depend continuously on $x$, such that $T_{x}M=H_{x}\oplus V_{x}$ and for the sets
$K^{s}(x):=\{(v_{1},v_{2})\in H_{x}\oplus V_{x}:|v_{1}|\leq c|v_{2}|\},$
$K^{u}(x):=\{(v_{1},v_{2})\in H_{x}\oplus V_{x}:|v_{2}|\leq c|v_{1}|\},$
we have
  • ${D_{x}f\left(K^{u}(x)\right)}\subseteq\text{int}K^{u}(f(x))$ and ${D_{x}f^{-1}\left(K^{s}(x)\right)}\subseteq\text{int}K^{s}\left(f^{-1}(x)\right)$
  • $\left\|D_{x}fv\right\|\geq\mu\|v\|$ for $v\in K^{s}(x)$
  • $\left\|D_{x}f^{-1}v\right\|\geq\mu\|v\|$ for $v\in K^{u}(\boldsymbol{x})$
[…]
It is well-known from [HP] that the stable and unstable sets of $\boldsymbol{x}\in M$ given by
$W^{s}(\boldsymbol{x}):=\left\{\boldsymbol{y}\in M:\lim\limits_{n\rightarrow\infty}\left|f^{n}(\boldsymbol{x})-f^{n}(\boldsymbol{y})\right|=0\right\}$
and
$W^{u}(\boldsymbol{x}):=\left\{\boldsymbol{y}\in M:\lim\limits_{n\rightarrow\infty}\left|f^{-n}(\boldsymbol{x})-f^{-n}(\boldsymbol{y})\right|=0\right\}$
are images of $C^{r}$ injective immersions from $\mathbb{R}$ to $M$. Hence, it is appropriate to refer to these sets as stable, respectively unstable, manifolds. When $\Lambda=M$, and since $\text{dim}(M)=2$, these objects can be used to construct Markov partitions of $M$ that are composed of Markov rectangles of arbitrarily small diameter (see Appendix 2 of [PT]).
If $\overline{f}=(f_{n})$ is a sequence of $C^{1}$ diffeomorphisms of $M$, then we say $\overline{f}$ satisfies a common cone condition on $M$ if for each $x\in M$, there are cones $K^{s}(x)$ and $K^{u}(x)$ as above such that for each $n\in\mathbb{N}$, we have
  • ${D_{x}f_{n}\left(K^{u}(x)\right)}\subseteq\text{int}K^{u}(f_{n}(x))$ and ${D_{x}f^{-1}_{n}\left(K^{s}(x)\right)}\subseteq\text{int}K^{s}\left(f^{-1}_{n}(x)\right)$
  • $\left\|D_{x}f_{n}v\right\|\geq\mu\|v\|$ for $v\in K^{s}(x)$
  • $\left\|D_{x}f^{-1}_{n}\right\|\geq\mu\|v\|$ for $v\in K^{u}(x)$
[…]
$W^{s}(\boldsymbol{x},m):=\left\{\boldsymbol{y}\in M:\lim\limits_{n\rightarrow\infty}\left|f_{n+m}\circ\cdots\circ f_{m}(\boldsymbol{x})-f_{n+m}\circ\cdots\circ f_{m}(\boldsymbol{y})\right|=0\right\}$
is the image of an injective $C^{r}$ immersion from $\mathbb{R}$ to $M$.
This result can be taken from the theory of random hyperbolic dynamics (see Chapter 7 of [Ar] and references therein), or one can use arguments similar to Theorem 6.2.8 from [KH] in combonation with Proposition 1, which involves graph transform techniques, to obtain this result. More general versions of this statement for Anosov families have been studied in [Mu]. We use a version of this statement, and sketch the proof, for our relevant family of hyperbolic maps in Section 3.4.
1.1.4 Thickness of Cantor Sets
A fundamental tool we will use to estimate Hausdorff dimension of Cantor sets is the notion of thickness for Cantor sets. Given a Cantor set $\mathcal{C}\subset\mathbb{R}$ with convex hull $I$, we say that $U\subset I\setminus\mathcal{C}$ is a gap of $\mathcal{C}$ if it is a connected component of $I\setminus\mathcal{C}$. Any enumeration $\mathcal{U}=\{U_{n}\}$ of the gaps of $\mathcal{C}$ is called a presentation of $\mathcal{C}$. Given a boundary point $u\in\mathcal{C}$ of some gap $U_{k}$, we call the connected component $B$ contained in $I\setminus\left(U_{1}\cup\dots\cup U_{k}\right)$ that contains $u$ a bridge of $U_{k}$. We define
$\tau(\mathcal{C},\mathcal{U},u):=\frac{|B|}{|U_{k}|},$
and set the thickness of $\mathcal{C}$ to be
(see Chapter 4 of [PT]). It is clear that these notions carry over to a Cantor set that belongs to a $C^{1}$ curve $\gamma\subset S^{1}$ where the gaps and bridges are measured using length relative to the curve $\gamma$.
1.1.5 Continued Fraction Expansion
The approach to our proof begins with some observations about irrational rotations on $S^{1}$ and their connections to Diophantine approximations.
$\frac{p_{n}(\alpha)}{q_{n}(\alpha)}=[a_{1},a_{2},\dots,a_{n}].$
$p_{k+1}(\alpha)=a_{k+1}p_{k}(\alpha)+p_{k-1}(\alpha),\ p_{1}(\alpha)=p_{0}(\alpha)=1$ (1.4)
$q_{k+1}(\alpha)=a_{k+1}q_{k}(\alpha)+q_{k-1}(\alpha),\ q_{1}(\alpha)=a_{1},\ q_{0}(\alpha)=1$ (1.5)
and that
$\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|<\frac{1}{q_{k}(\alpha)}.$ (1.6)
and by (1.6), both of these quantities are no greater than $\frac{1}{q_{k-1}(\alpha)}+\frac{1}{q_{k}(\alpha)}$. As a corollary, we obtain the following result which will be used later on.
Proposition 2.
$\min\limits_{1\leq m\leq n}\left|m\alpha-y\right|<\epsilon$
for any $y\in S^{1}$.
Proof.
Let $\epsilon>0.$ Set $Y:=\{\alpha=[a_{1},a_{2},\dots]:a_{i}\in\mathcal{A}\}\subset[0,1]\setminus\mathbb{Q}$ and let $a^{\prime}$ be the smallest element of $\mathcal{A}$. Then using the recurrence relation (1.5) and an induction argument, we find that for each $k\geq 1$, there is an $M_{k}>0$ such that
Choose $k$ large enough so that
1.2 Outline of Paper
Let us fix an irrational $\alpha=[a_{1},a_{2},\dots]\in(0,1)$ such that the sequence $(a_{n})_{n\in\mathbb{N}}$ is bounded. Since we are working in a small range of $\lambda$, our hope is to compare the dynamics of our trace maps restricted to $S_{\lambda}$, to the dynamics of the trace maps on $S_{0}$. However, this comparison becomes difficult near the cusps of $S_{\lambda}$ due to how the curvature of the surface changes as $\lambda\rightarrow 0$.
@@ 1.2 Outline of Paper
For any points that has a bounded orbit, its orbit may enter a neighborhood of one of the cusps of $S_{\lambda}$ for some amount of time but then return to a central portion of the surface where the dynamics are comparable to those on $S_{0}$. This behavior results in the need for technical distortion estimates, and makes obtaining a common cone condition for our family of trace maps, over these cusps, rather difficult. These details have been written in Section 3 of [DG3] for the map $T_{1}$ only.
@@
To avoid this obstacle, we will first remove neighborhoods of $S_{\lambda}$ near the cusps, so that the remaining surface is a union of five disjoint components, only one of which is bounded. We refer to this bounded component as a central portion of $S_{\lambda}$. We then turn our attention to the set of points from this central portion, whose orbits never leave this region. The points from $L_{\lambda}$ that satisfy this behavior form a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$, and it is sufficient to show that this subset has Hausdorff dimension near $1$ as $\lambda$ approaches 00. Provided $\lambda$ is sufficiently small, it turns that each trace map restricted to a desired central portion of $S_{\lambda}$ is a factor of an Anosov map restricted to $\mathbb{T}^{2}$ with a small open set removed, and this map is a perturbation of a hyperbolic linear automorphism.
@@
  • In the beginning of Section 3, using some results from 2, we find an open set $O\subset\mathbb{T}^{2}$ composed of four small neighborhoods, a sequence of maps $\left(\tilde{T}_{\lambda,a_{n}}\right)$ defined on $\mathbb{T}^{2}\setminus O$, and a curve $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus O$ that tends to the line $\left[0,\frac{1}{2}\right]\times\{0\}$ as $\lambda\rightarrow 0$, such that the set $\mathcal{C}({\lambda},O):=\left\{P\in\hat{L}_{\lambda}:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O\ \text{for all}\ n>0\right\}$ is a Cantor set in $\hat{L}_{\lambda}$ that has dimension equal to the dimension of a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$ for a small range of $\lambda$. The rest of the paper is dedicated to showing that for this small range of $\lambda$, the Hausdorff dimension of $C(\lambda,O)$ is close to $1$.
  • Our approach to estimating the dimension of $C(\lambda,O)$ is to show that for small $\lambda$, there is a Cantor set of large thickness that is contained in $C(\lambda,O)$. This is the goal of Section 3, and to do so, we extend each map $\tilde{T}_{\lambda,a_{n}}$ to an Anosov diffeomorphisms on $\mathbb{T}^{2}$, and analyze the non-stationary dynamics of these extended maps. This Anosov family will be constructed to satisfy a common cone condition, and we derive a non-stationary version of the Stable Manifold Theorem for these maps in Section 3.4.
  • Section 3.5 is dedicated to constructing a particular set of dynamical rectangles, which will help us estimate the dimension of $\mathcal{C}(\lambda,O)$. Roughly speaking, we will construct a family of curvilinear rectangles that have two opposite boundaries composed of pieces of the stable manifolds developed in Section 3.4, and these rectangles will have controlled size and contain the open set $O$. We use the common cone conditions to control the size of these rectangles.
  • In Section 3.6, we use the properties of the dynamical rectangles constructed in Section 3.5 to construct a Cantor set of large thickness that is contained in $C(\lambda,O)$. Roughly speaking, this Cantor set is derived from considering the points of $\hat{L}_{\lambda}$ that do not enter a family of associated rectangles containing $O$, under our non-stationary iterations of extended Anosov maps. The dynamical nature of these rectangles allows us to estimate the thickness of the newly derived Cantor set. We use this result to prove Theorem 1.
As a choice of notation, we will always use dist to denote the Hausdorff distance between sets and $d_{C^{0}}$ to denote the $C^{0}$ distance between mappings. With a slight abuse of notation, we will denote $(x,y)$ to be the standard coordinates in both $\mathbb{T}^{2}$ and $\mathbb{R}^{2}$.

### v2

@@ 1 Introduction
This operator falls under the category of a discrete Schrödinger operator with dynamically defined potential. This so-called Sturmian potential has been studied since the early 1980s, and has been a popular model in the study of one-dimensional quasicrystals. A Sturmian sequence is an aperiodic sequence of least possible complexity, and hence relates to the structure of quasicrystals which are highly ordered and aperiodic. A survey of results concerning these types of operators, especially Schrödinger operators, can be found in [DEG, DF2].
We focus our attention to the case when the frequency has only finitely many different terms that appear in its continued fraction expansion (i.e. of bounded-type) and the coupling is weak (i.e. $\lambda$ is near 00), and investigate the dimension of the spectrum of $H_{\lambda,\alpha,\omega}$ in this context. Since $R_{\alpha}$ is minimal, it is well-known (see Theorem 4.9 in [DF1]) that the spectrum of $H_{\lambda,\alpha,\omega}$ is independent of $\omega$, so we will denote it by $\sigma_{\lambda,\alpha}$.
@@
In the case that $\alpha=\frac{\sqrt{5}-1}{2}$, the operator $H_{\lambda,\alpha,\omega}$ is known as the Fibonacci Hamiltonian. It was first discovered that the spectrum of this operator is a Cantor set of zero Lebesgue measure in [Su2] and many questions regarding the box-counting and Hausdorff dimensions of the spectrum have been answered (see [D1, D2, DEG, DGY, Su3] for surveys).
@@
For weak coupling constant, many dimension results about the spectrum have been obtained through the analysis of trace map formalism. There is a fundamental link between the spectrum $\sigma_{\lambda,\alpha}$ and the dynamics of the so-called trace maps over a family of cubic surfaces. More precisely, for each $a\in\mathbb{N}$, there is a corresponding trace map $T_{a}$ which leaves invariant a corresponding family of cubic surfaces $\{S_{\lambda}\}_{\lambda>0}$, and the dynamics of the trace maps over $S_{\lambda}$ in the prescribed order that comes exactly from the continued fraction expansion of $\alpha$, gives rich information about the spectrum $\sigma_{\lambda,\alpha}$ (see [Su1, Ra]). For example, the continued fraction expansion of $\alpha=\frac{\sqrt{5}-1}{2}$ contains only 1’s, so here, only the Fibonacci trace map $T_{1}$ plays an important role. Many authors have examined dynamical properties of this trace map (see [BGJ, BR, C, Ca, DG1, DG2, DG3, DG4, DG5, KKT]), where questions concerning hyperbolicity, Markov partitions, entropy, and more have been addressed. It turns out that the set of points with bounded orbit under the Fibonacci trace map restricted to any $S_{\lambda}$, $\lambda>0$, is a locally maximal hyperbolic set that is homeomorphic to a Cantor set. This was proved for $\lambda\geq 16$ in [C], for small $\lambda$ in [DG1], and for all $\lambda$ in [Ca]. When $\alpha$ is eventually periodic, one needs to investigate the non-wandering set of a hyperbolic map over $S_{\lambda}$, where this map is precisely a composition of finitely many trace maps, in which case several well-developed techniques from hyperbolic dynamics apply.
In this paper, to work with the weakened assumption that $\alpha$ is of bounded-type, we incorporate non-stationary versions of various dynamical notions such as stable manifolds, Markov rectangles, and a bounded distortion estimate from [BM], since many of these classical results cannot be directly applied.
@@ 1.1 Background Material
A main ingredient for achieving such results about the spectrum $\sigma_{\lambda,\alpha}$ is to investigate the dynamics of the so-called trace maps. Consider the maps $G,H:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ via
$G(x_{1},x_{2},x_{3}):=(2x_{1}x_{3}-x_{2},x_{1},x_{3})\ \text{and}\ H(x_{1},x_{2},x_{3}):=(x_{1},x_{3},x_{2}).$
For $a\in\mathbb{N}$, we define the $a^{\text{{th}}}$-trace map to be $T_{a}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ via
$T_{a}=G^{a}\circ H.$ (1.2)
@@
Each surface $S_{\lambda}$ is invariant under each trace map. Given a sequence $\overline{a}=(a_{n})$, $a_{n}\in\mathbb{N}$, denote
$\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right):=\left\{\boldsymbol{x}\in S_{\lambda}:\left(T_{a_{n}}\circ\cdots{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ}T_{a_{1}}(\boldsymbol{x})\right)_{n\in\mathbb{N}}\ \text{is bounded}\right\}.$
@@
which is contained in $S_{\lambda}$, then our goal is to estimate the Hausdorff dimension of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$ as $\lambda\rightarrow 0$ (see Figure 1).
Figure 1: Surface $S_{\frac{1}{4}}$ and line $L_{\frac{1}{4}}$.
1.2 Approach and Outline of Paper
Consider the points $P_{1}=(1,1,1),\ P_{2}=(-1,-1,1),\ P_{3}=(1,-1,-1),$ and $P_{4}=(-1,1,-1)$. For $\rho>0$, let us set $O_{\rho}:=\bigcup_{i=1}^{4}B_{\rho}\left(P_{i}\right),$ so that for small enough $\lambda$, we will have that $S_{\lambda}\setminus O_{\rho}$ is composed of five disjoint components, only one of which is bounded (see Figure 2). Denote this bounded component by $\mathbb{S}_{\lambda,O_{\rho}}$. For a fixed sequence $\overline{a}=(a_{n})$, $a_{n}\in\mathbb{N},$ and for small enough $\lambda$, we define
@@ 1.2 Approach and Outline of Paper
$\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right):=\left\{\boldsymbol{x}\in\mathbb{S}_{\lambda,O_{\rho}}:T_{\lambda,a_{n}}\circ\cdots\circ T_{\lambda,a_{1}}(\boldsymbol{x})\in\mathbb{S}_{\lambda,O_{\rho}},\ \text{for all}\ n\in\mathbb{N}\right\}.$
This set can be thought of as a non-stationary survival set. We will then prove
Theorem 3.
Let $\alpha=[a_{1},a_{2},\dots]\in(0,1)$ be an irrational of bounded type and set $\overline{a}=(a_{n})$. For each $\epsilon>0$, there is a $\rho>0$ and $\lambda_{0}>0$ such that
$\text{dim}_{H}\left(L_{\lambda}\cap\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)\right)\geq 1-\epsilon$
for all $\lambda\in[0,\lambda_{0}]$.
Since $\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)\subset\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$, together with Theorem 2, this will immediately imply Theorem 1.
Figure 2: Depiction of $S_{\lambda}\setminus O_{\rho}$.
The motivation for studying the set $\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)$ comes from the fact that near each $P_{i}$, the curvature of the surface $S_{\lambda}$ becomes large as $\lambda\rightarrow 0$. This makes analyzing the trace map dynamics near these cusps complicated.
@@
For any point that has a bounded orbit, its orbit may enter a neighborhood of one of the cusps of $S_{\lambda}$ for some amount of time but then return to a central portion of the surface where the dynamics are comparable to those on $S_{0}$. This behavior results in the need for technical distortion estimates, and makes obtaining a common cone condition for our family of trace maps, over these cusps, rather difficult. These details have been written in Section 3 of [DG3] for the map $T_{1}$ only.
@@
So, to avoid this obstacle, we will first remove neighborhoods of $S_{\lambda}$ near the cusps, and instead study the set of points whose orbits remain (or survive) in the left over bounded component. The points from $L_{\lambda}$ that satisfy this behavior form a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$, and it is sufficient to show that this subset has Hausdorff dimension near $1$ as $\lambda$ approaches 00. Provided $\lambda$ is sufficiently small, it turns out that each trace map restricted to a desired central portion of $S_{\lambda}$ is a factor of an Anosov map restricted to the two torus $\mathbb{T}^{2}:=\mathbb{R}^{2}/\mathbb{Z}^{2}$ with a small open set removed, and this map is a perturbation of a hyperbolic linear automorphism. Thus, as we will see later in the paper (see Section 4), studying the set $\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)$ can be translated to a problem involving a sequence of toral Anosov maps.
@@
  • In Section 4, we state and prove Theorem 4, which is a translation of Theorem 3 to a problem concerning the non-stationary dynamics of a sequence of toral maps. The proof of this statement is done through the entirety of Section 4, and hence, we give an outline of the argument immediately after the statement of Theorem 4.
  • In Section 5, we use Theorem 4 to conclude the proofs of Theorem 3 and Theorem 1.
As a choice of notation, we will always use dist to denote the minimal distance between sets and $\text{dist}_{H}$ to denote the Hausdorff distance between sets and $d_{C^{0}}$ to denote the $C^{0}$ distance between mappings. With a slight abuse of notation, we will denote $(x,y)$ to be the standard coordinates in both $\mathbb{T}^{2}$ and $\mathbb{R}^{2}$.